OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP!!! Use the one to one property to solve the given exponential equation... 3^x * 5^x = 225

OpenStudy (freckles):

try to find prime factorization for 225

OpenStudy (freckles):

for example" \[3^x5^x=15 \\ \text{ prime factorization for } 15 \text{ is } 3 \cdot 5 \\ 3^x 5^x=3 \cdot 5 \\ 3^x5^x=3^1 \cdot 5^1 \\ \text{ so } x=1 \]